# Socialite-llama: An Instruction-Tuned Model for Social Scientific Tasks

### 03 Feb 2024 v1

@@ 3.1 Training Tasks
Sentiment and Emotion For emotion classification, we use SemEval 2018 Task 1 data Mohammad et al. (2018). Emotion involves classifying a tweet into anger, joy, optimism or sadness displayed by the author. We also include Emobank Buechel and Hahn (2017) to infer the valence, arousal and dominance levels that a text would invoke in a reader. We convert these regression tasks to binary classification - by transforming a score greater than 4 to ’high’ and less than 4 to ’low’, for each of the three datasets respectively (ValenceCLS, ArousalCLS, DominanceCLS). We use SemEval 2017 Task 4 data for sentiment classification Rosenthal et al. (2017) to identify the overall sentiment of a text as positive, negative or neutral. Further, we also include SameSideStance Körner et al. (2021), the task of classifying whether two different texts are on the same side of an argument.
@@ Acknowledgements
We would like to thank the reviewers for their valuable feedback that helped us improving this paper. This work was supported in part by a grant from the CDC/NIOSH (U01 OH012476), a grant from IARPA (585968), a grant from the NIH-NIAAA (R01 AA028032). YKL was supported in part by the Air Force Research Laboratory (AFRL), DARPA, for the KAIROS program under agreement number FA8750-19-2-1003 and in part by the NSF under the award IIS #2007290. We approve this work for public release and redistribution with attribution. The conclusions and views in this paper are those of the authors, and should not be seen as representing the official policies expressed or implied by CDC, NIOSH, NIH, NIAAA, IARPA, DARPA, ARFL, NSF or the U.S. government.
@@ Appendix A Appendix
Instruction: Analyze the provided text using the Valence-Arousal-Dominance (VAD) emotional model. Your task is to classify the arousal level it might trigger in an average reader. Arousal, in this context, refers to the degree of energy or lethargy the text might induce. ’Low Arousal’ indicates a low arousal level, suggesting the text is likely to make the reader feel calm or lethargic. Conversely, ’High Arousal’ indicates a high arousal level, suggesting the tweet is likely to energize or excite the reader. Your options for classification are confined to ‘Low Arousal’ or ‘High Arousal’.
@@
Instruction: Analyze the provided text, considering its tone and language, and categorize it as either polite or impolite. Your options for classification are confined to ’impolite’ or ’polite’.

### 14 Mar 2024 v2

@@ 3.1 Training Tasks
Sentiment and Emotion For emotion classification, we use SemEval 2018 Task 1 data Mohammad et al. (2018). Emotion involves classifying a tweet into anger, joy, optimism or sadness displayed by the author. We also include Emobank Buechel and Hahn (2017) to infer the valence, arousal and dominance levels that a text would invoke in a reader. We convert these regression tasks to binary classification - by transforming a score greater than 4 to ‘high’ and less than 4 to ‘low’, for each of the three datasets respectively (ValenceCLS, ArousalCLS, DominanceCLS). We use SemEval 2017 Task 4 data for sentiment classification Rosenthal et al. (2017) to identify the overall sentiment of a text as positive, negative or neutral. Further, we also include SameSideStance Körner et al. (2021), the task of classifying whether two different texts are on the same side of an argument.
@@ Acknowledgements
We would like to thank the reviewers for their valuable feedback that helped us improving this paper. This work was supported in part by the Office of the Director of National Intelligence (ODNI), Intelligence Advanced Research Projects Activity (IARPA), via the HIATUS Program contract (585968), a grant from the CDC/NIOSH (U01 OH012476), and a grant from the NIH-NIAAA (R01 AA028032). YKL was supported in part by the Air Force Research Laboratory (AFRL), DARPA, for the KAIROS program under agreement number FA8750-19-2-1003 and in part by the NSF under the award IIS 2007290. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of ODNI, IARPA, any other government organization, or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright annotation therein.
@@ Appendix A Appendix
Instruction: Analyze the provided text using the Valence-Arousal-Dominance (VAD) emotional model. Your task is to classify the arousal level it might trigger in an average reader. Arousal, in this context, refers to the degree of energy or lethargy the text might induce. ‘Low Arousal’ indicates a low arousal level, suggesting the text is likely to make the reader feel calm or lethargic. Conversely, ‘High Arousal’ indicates a high arousal level, suggesting the tweet is likely to energize or excite the reader. Your options for classification are confined to ‘Low Arousal’ or ‘High Arousal’.
@@
Instruction: Analyze the provided text, considering its tone and language, and categorize it as either polite or impolite. Your options for classification are confined to ‘impolite’ or ‘polite’.